OpenStudy (sbentleyaz):

The side of a square is measured to be 12 ft with a possible error of ±0.1 ft. Use differentials to estimate the error in the calculated area. Include units in your answer

OpenStudy (mathmale):

Write the Area function: A(x)=x^2. Differentiate this with respect to x. Multiply both sides of the resulting equation by dx. Your result? Now replace x in your result with 12 ft, and replace dx with (plus or minus) 0.1 ft.

OpenStudy (mathmale):

The resulting equation, for dA, represents the approx error in the area. Topic: "Differentials."

OpenStudy (sbentleyaz):

so its A(x)=2x?

OpenStudy (sbentleyaz):

then d(x)A(x)=2x(d(x))

OpenStudy (mathmale):

Actually, you want A '(x) = 2x dx, or\[\frac{ dA }{ dx }=2x\] Multiply both sides of this equation by dx and simplify the result.

OpenStudy (sbentleyaz):

so dxDa/dx=2xdx dA=2xdx dA=\[24\pm.1\]

OpenStudy (sbentleyaz):

is that correct?

OpenStudy (sbentleyaz):

@TheSmartOne

OpenStudy (sbentleyaz):

@3mar

OpenStudy (irishboy123):

this, in differential language, is: \(A = x^2\) \(dA = 2x ~ dx\) A very cheap but ubiquitous trick is this - take logs of \(A = x^2\): \(\implies \ln A = \ln x^2 = 2 \ln x\) then \(d ( \ln A = 2 \ln x) \implies \dfrac{dA}{A} = 2 \dfrac{dx}{x} \implies dA =2A \dfrac{dx}{x}\) if \(A = x^2\) then you get the same result. but really it's all about Taylor Expansions \(A(x + \delta x) = A(x) +\delta x A'(x) + \mathcal O (\delta x^2)\) \(\delta A = A(x + \delta x) - A(x) = \delta x f'(x) + \mathcal O (\delta x)\) \(\implies \delta A = \delta x (2x) + \mathcal O (\delta x)\) \(\implies \delta A \approx (\pm 1) . (24) \)

OpenStudy (sbentleyaz):

it seems really unlikely the error would be 14.4 isn't that a really high number

OpenStudy (irishboy123):

typo \(\delta A \approx (\pm \color{red}{0.}1) . (24) = \pm 2.4\)

OpenStudy (sbentleyaz):

so would the final answer be plus or minus 2.4 feet squared @IrishBoy123

OpenStudy (mathmale):

144 square feet plus or minus 2.4 square feet.